Calculate as x approaches -8(+)(right hand limit) of 2x/x+8.

Grade:12

2 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
L = \lim_{x\rightarrow -8^{+}}\frac{2x}{x+8}
-8\rightarrow -8+h, h>0
L = \lim_{h\rightarrow 0}\frac{2(-8+h)}{-8+h+8}
L = \lim_{h\rightarrow 0}\frac{-16+2h}{h} = +\infty
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty
Jitender Singh
9 years ago
Sorry about the previous answers, there is an minor mistake.
L = (-ve) Infinity
 
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty